# Remasking Discrete Diffusion Models with Inference-Time Scaling

### v1

@@ 1 Introduction
Diffusion models have gained significant traction as algorithms for generating high-quality images and videos (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020). Part of the success of diffusion stems from its ability to perform iterative refinement—repeatedly modifying outputs and fixing errors over multiple steps of generation—which makes diffusion models especially effective at guided generation (Dhariwal & Nichol, 2021; Ho & Salimans, 2022), fast sampling (Song et al., 2020a; Salimans & Ho, 2022; Song et al., 2023), and supports inference-time scaling to improve sample quality (Ma et al., 2025).
Discrete counterparts of diffusion models (Austin et al., 2021) have also been steadily improving in quality on tasks such as language modeling and biological sequence design (Lou et al., 2024; Sahoo et al., 2024; Schiff et al., 2024). However, modern discrete diffusion—especially the most performant kind that relies on masking (Ou et al., 2024; Shi et al., 2024; Sahoo et al., 2024)—lacks the fundamental ability to iteratively refine outputs. Once a discrete token is unmasked, it cannot be updated again, even if it introduces an error. The lack of error correction in turn sets limits on controllable generation, sampling speed, and sample quality.
@@
Figure 1: Our family of masked diffusion processes allow for more flexible generation with remasking of already decoded tokens. This improves sample quality and further closes the gap to AR models. (Left) An illustrative example of errors fixed by ReMDM. The first two tokens can be “They sell” or “She sells”, but due to the independence of the parallelized decoding processes, “She sell” is decoded. Such mistakes can be corrected by remasking samplers. (Right) MAUVE scores on OpenWebText. MDLM is from Sahoo et al. (2024). FB and DFM denote the forward-backward (Campbell et al., 2022) and discrete flow matching (Gat et al., 2024) correctors, respectively.
Our approach for deriving the sampler is rooted in probabilistic modeling—-we show that our method corresponds to ancestral sampling in a discrete diffusion model (also called ReMDM) characterized by a remasking backward process. The ReMDM model admits an objective that is a rescaled version of the MDLM objective (Sahoo et al., 2024)—this suggests using the ReMDM sampler on top of pre-trained MDLMs, which we find to work well. We complement our analysis of the sampler by also showing that it can be interpreted as a predictor-corrector technique (Song et al., 2020b; Campbell et al., 2022).
Most interestingly, the ReMDM sampler endows discrete diffusion with a form of inference-time compute scaling. By increasing the number of denoising steps on language modeling tasks, ReMDM generates samples with significantly higher sample quality metrics than any previous diffusion model and almost matches the performance of an autoregressive (AR) model with the same architecture. Conversely, when we reduce sampling steps to increase speed, performance degrades less than other samplers on text and image generation. Lastly, ReMDM models are more amenable to guidance (Schiff et al., 2024)—in molecule design experiments, ReMDM pushes the Pareto frontier of novelty and the desired molecular property relative to alternatives relying on masking or uniform noise diffusion.
@@
    We introduce the remasking diffusion model (ReMDM) sampler and a number of add-on components that bring performant iterative refinement via remasking to masked diffusion models.
@@
    We show that our method is a form of ancestral sampling in a probabilistic model whose ELBO is similar to that of classical masked diffusion. This analysis suggests using our sampler on top of pre-trained models, which we find to work well.
@@
    Across the domains of natural language, discretized images, and molecule string representations, we demonstrate empirically that ReMDM endows masked diffusion with inference-time scaling that improves sample quality with more computation, and that also enhances controlled generation.
@@ 2 Background
Discrete Diffusion Models  Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are characterized by parametric models $p_{\theta}$ trained to reverse a fixed noising process $q$ that maps clean data $\mathbf{x}$ to increasingly noisy latent variables $\mathbf{z}_{t}$, for $t\in[0,1]$. Austin et al. (2021); Sahoo et al. (2024); Shi et al. (2024) extend this framework to discrete signals. Formally, we define $\mathbf{x}\in\{0,1\}^{|V|}\subset\Delta^{|V|}$ as one-hot vectors, where $|V|$ is the vocabulary size and $\Delta^{|V|}$ is the simplex over this vocabulary. For finite-time processes, we let $T$ be the number of time steps. For each time step $t(i)=\frac{i}{T}\in[0,1]$ (i.e., for $i=0,1,\ldots,T$), each intermediate latent variable $\mathbf{z}_{t(i)}\in\{0,1\}^{|V|}$ has the marginal:
@@
Masked Diffusion Language Models  A special case of this formulation, known as ‘masking’ or ‘absorbing state’ discrete diffusion, sets the limiting distribution $\bm{\pi}=\bm{m},$ a one-hot vector centered on a special [MASK] token. Letting $s(i)=\frac{i-1}{T}$ be the time step directly preceding $t$, these posteriors take the form:
@@
Importantly, despite the recent success of MDLM (Sahoo et al., 2024) and other absorbing state discrete diffusion models, these models suffer from a key drawback, which we call the failure to remask property. Namely, the posterior formula in (2) implies that any unmasked token $\mathbf{z}_{t}\neq\bm{m}$ must remain unchanged throughout the entire denoising process, for all $s\leq t,$ which Sahoo et al. (2024) formalize as $q(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})=\mathrm{Cat}(\mathbf{z}_{s};\mathbf{z}_{t})$ if $\mathbf{z}_{t}\neq\bm{m}$. Once a token is decoded, this prediction is ‘locked-in’, a limitation that these diffusion models share with AR models.
As in Sahoo et al. (2024), a denoising neural network $\mathbf{x}_{\theta}$ is used for parameterizing $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})=q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{\theta}(\mathbf{z}_{t}))$ and trained using the following objective:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}_{0}\sim q(\mathbf{z}_{0}\mid\mathbf{x})}\bigg{[}-\log p_{\theta}(\mathbf{x}\mid\mathbf{z}_{0})\bigg{]}$ (3)
$+\mathbb{E}_{t\in\{\frac{1}{T},\ldots,1\}}\mathbb{E}_{\mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{x})}T\bigg{[}\frac{\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\log(\mathbf{x}_{\theta}(\mathbf{z}_{t})^{\top}\mathbf{x})\bigg{]}$
Discrete Guidance  An important property of diffusion models is their amenability to controlled generation through the mechanisms of classifier-based (CBG; Dhariwal & Nichol (2021)) and classifier-free guidance (CFG; Nichol et al. (2021); Ho & Salimans (2022)). Recently, these mechanisms have been extended to discrete diffusion (Nisonoff et al., 2024; Schiff et al., 2024). Letting, $\gamma$ represent a hyperparameter that controls guidance strength and $y$ the label of the desired class output, following the presentation in Schiff et al. (2024), we implement discrete CBG (D-CBG) by drawing samples for $\mathbf{z}_{s}^{(\ell)}$ from a tempered distribution:
$p^{\gamma}(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)},y)\propto p_{\phi}(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)})^{\gamma}\cdot p_{\theta}(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}),$ (4)
where $p_{\phi}$ is a classifier trained on noisy sequences and evaluated on sequences where the $\ell^{\text{th}}$ token in $\mathbf{z}_{t}^{(1:L)}$ is replaced by $\mathbf{z}_{s}^{(\ell)}$. For discrete CFG (D-CFG), we sample from
$\begin{split}&p^{\gamma}(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)},y)\propto\\ &~{}~{}~{}~{}~{}~{}~{}~{}p_{\theta}(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)},y)^{\gamma}\cdot p_{\theta}(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)})^{(1-\gamma)},\end{split}$ (5)
where the same denoising network $\mathbf{x}_{\theta}$ is used to model both the conditional and unconditional distributions, with the class label simply masked-out for the latter.
@@ 3 Remasking Diffusion Models
In this work, we alleviate the failure to remask limitation that affects state-of-the-art masked diffusion models. Our strategy is to design a probabilistic model similar to MDLM (Sahoo et al., 2024), but whose denoising posterior generalizes (2) and supports remasking.
Specifically, we will define a discrete diffusion model whose posterior for $\mathbf{z}_{t}\neq\bm{m}$ has the following form:
@@ 3.1 ReMDM Forward Processes
where $q_{\sigma}(\mathbf{z}_{1}\mid\mathbf{x})=q(\mathbf{z}_{1})=\bm{m}$ is the same limiting distribution as in masked diffusion. We construct the posteriors $q_{\sigma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})$ as:
$\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t})\mathbf{x}+\sigma_{t}\bm{m}),\quad\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}+\frac{1-\alpha_{s}-\sigma_{t}\alpha_{t}}{1-\alpha_{t}}\bm{m}),\quad\mathbf{z}_{t}=\bm{m}.\end{cases}$ (8)
Observe how (8) maintains the remasking property laid out in (6). Additionally, our chosen form for $q(\mathbf{z}_{s}\mid\mathbf{z}_{t}=\mathbf{m},\mathbf{x})$ ensures that the marginals for $q_{\sigma}(\mathbf{z}_{t}\mid\mathbf{x})$ are the same as in classical masked diffusion (e.g., Sahoo et al. (2024)). This property will yield an ELBO for ReMDM that is very similar to that of MDLM, and will help us argue for using the ReMDM sampling procedure on top of a pre-trained MDLM model. Formally, we establish the following result (proof in Appendix A.1):
@@
In Appendix A.2, we demonstrate that this ability to remask necessitates that ReMDM is a non-Markovian process (as in e.g., DDIM (Song et al., 2020a)). Intuitively, this is necessary because the latents $\mathbf{z}_{t}$ can transition from the masked state $\bm{m}$ back to $\mathbf{x}$, and therefore the transition kernel needs to be conditioned on $\mathbf{x}$.
@@ 3.2 Negative Evidence Lower Bound
$\mathcal{L}^{\sigma}=\mathbb{E}_{\mathbf{z}_{0}\sim q(\mathbf{z}_{0}\mid\mathbf{x})}\bigg{[}-\log p_{\theta}(\mathbf{x}\mid\mathbf{z}_{0})\bigg{]}$ (11)
$+$ $\mathbb{E}_{t\in\{\frac{1}{T},\ldots,1\}}\mathbb{E}_{\mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{x})}T\bigg{[}\frac{(1-\sigma_{t})\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\log(\mathbf{x}_{\theta}^{\top}\mathbf{x})\bigg{]}.$
Since $\alpha_{t}\leq\alpha_{s},$ each term in the second expectation increases monotonically in $\sigma_{t}$. When $\sigma_{t}=0$, for all $t$, we recover the MDLM objective from (3). This implies that the MDLM NELBO produces a tighter bound compared to ReMDM. Furthermore, since the diffusion loss term in (11) is simply a reweighted version of that in (3), one can presumably reuse weights from a model trained with (3), i.e., using different $\sigma_{t}$ at training and inference. Indeed, we find this to be a performant strategy in practice. Moreover, we find the performance (test perplexity) between models trained with the MDLM and the ReMDM objectives to be comparable (see Appendix E.1), further justifying our reuse of pre-trained weights, with the added benefit of more flexible sampling unlocked by ReMDM.
@@ 4 ReMDM Samplers
In practice, we recommend using ReMDM with a different $\sigma_{t}$ at training and inference, just as one might switch to a different noise schedule during sampling. In particular, using $\sigma_{t}=0$ for training is equivalent to combining the ReMDM sampler with a pre-trained MDLM model $\mathbf{x}_{\theta}$, which works well and does not require re-training $\mathbf{x}_{\theta}$.
@@ 4.1 Design Strategies for the Remasking Schedule $\sigma_{t}$
Here, we list several strategies for setting $\sigma_{t}\in[0,\sigma_{t}^{max}]$, which can either be employed separately or in conjunction with the strategies introduced in Section 4.2. Note that for sequences, we can assign a different $\sigma_{t}^{(\ell)}$ for each token. When we omit this superscript, this implies that the same $\sigma_{t}$ is used for all tokens.
@@
Confidence-Based Schedule  In conjunction with the two strategies above, we explore a further reweighing of $\sigma_{t}$ which is based on the intuition that tokens of which the denoising model is less confident should be assigned a larger probability of remasking. Consider the $\ell$-th token in a sequence of $L$ latents at time $t$. For each $\ell\in\{1,\ldots,L\}$, we store its decoding probability at the time $\tau$ at which it was last unmasked. Concretely, if $\mathbf{z}_{t}^{(\ell)}\neq\bm{m}$, then we define $\psi_{t}^{(\ell)}:=\mathbf{x}_{\theta,\tau}^{(\ell)\top}\mathbf{z}_{\tau}^{(\ell)}$. If $\mathbf{z}_{t}=\bm{m}$, then $\psi_{t}^{(\ell)}:=\infty$. Thus, $\psi_{t}^{(\ell)}$ serves as a ‘confidence score’ for unmasked tokens.
We then compute
@@
$p_{\theta}(\mathbf{z}_{s}\mid\mathbf{z}_{t})=q_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\sigma}}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}=\mathbf{x}_{\theta}(\mathbf{z}_{t}))$
=$={\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t})\mathbf{x}_{\theta}+\sigma_{t}\bm{m}),&\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}_{\theta}+\frac{1-\alpha_{s}-\sigma_{t}\alpha_{t}}{1-\alpha_{t}}\bm{m}),&\mathbf{z}_{t}=\bm{m}\\ \end{cases}}$

### v2

@@ 1 Introduction
Diffusion models have gained significant traction as algorithms for generating high-quality images and videos [47, 49, 18, 39]. Part of the success of diffusion stems from its ability to perform iterative refinement—repeatedly modifying outputs and fixing errors over multiple steps of generation—which makes diffusion models especially effective at guided generation [9, 17] and fast sampling [48, 41, 51] and supports inference-time scaling to improve sample quality [27, 46, 61].
Discrete counterparts of diffusion models [2] have also been steadily improving in quality on tasks such as language modeling and biological sequence design [26, 38, 43, 1, 40]. However, modern discrete diffusion—especially the most performant kind that relies on masking [29, 38, 45]—lacks the fundamental ability to iteratively refine outputs. Once a discrete token is unmasked, it cannot be updated again, even if it introduces an error. The lack of error correction in turn sets limits on controllable generation, sampling speed, and sample quality.
@@
Figure 1: Our family of masked diffusion processes allow for more flexible generation with remasking of already decoded tokens. This improves sample quality and further closes the gap to AR models. (Left) An illustrative example of errors fixed by ReMDM. The first two tokens can be “They sell” or “She sells”, but due to the independence of the parallelized decoding processes, “She sell” is decoded. Such mistakes can be corrected by remasking samplers. (Right) MAUVE scores on OpenWebText. MDLM is from Sahoo et al. [38]. FB and DFM denote the forward-backward [5] and discrete flow matching [13] correctors, respectively.
Our approach for deriving the sampler is rooted in probabilistic modeling. We show that our method corresponds to ancestral sampling in a discrete diffusion model (also called ReMDM) characterized by a remasking backward process. The ReMDM model admits an objective that is a rescaled version of the MDLM objective [38]. This suggests using the ReMDM sampler on top of pretrained MDLMs, which we find to work well. We complement our analysis of the sampler by also showing that it can be interpreted as a predictor-corrector technique [50, 5]. Notably, we demonstrate that ReMDM is theoretically more general than previous corrector samplers for masked diffusion models [5, 13].
Most interestingly, the ReMDM sampler endows discrete diffusion with a form of inference-time compute scaling. By increasing the number of denoising steps on language modeling tasks, ReMDM generates samples with significantly higher sample quality metrics than any previous diffusion model and almost matches the performance of an autoregressive (AR) model with the same architecture. Conversely, when we reduce the sampling steps to increase speed, ReMDM’s performance degrades less than other samplers. On image generation, ReMDM  scales better with inference-time compute than vanilla masked diffusion models like MaskGiT [7]. Additionally, ReMDM models are more amenable to guidance [43]—in molecule design experiments, ReMDM pushes the Pareto frontier of novelty and the desired molecular property relative to alternatives relying on masking or uniform noise diffusion.
@@
    We introduce the remasking diffusion model (ReMDM) sampler that yields performant iterative refinement via remasking to masked diffusion models.
@@
    We show that our method is a form of ancestral sampling in a probabilistic model whose ELBO is similar to that of classical masked diffusion. This analysis suggests using our sampler on top of pretrained models, which we find to work well.
@@
    Across the domains of natural language, discretized images, and molecule string representations, we demonstrate empirically that ReMDM endows masked diffusion with inference-time scaling that improves sample quality with more computation and enhances controlled generation and downstream performance.
@@ 2 Background
Discrete Diffusion Models  Diffusion models [47, 18] are characterized by parametric models $p_{\theta}$ trained to reverse a fixed noising process $q$ that maps clean data $\mathbf{x}$ to increasingly noisy latent variables $\mathbf{z}_{t}$, for $t\in[0,1]$. Austin et al. [2], Sahoo et al. [38], Shi et al. [45] extend this framework to discrete signals. Formally, we define $\mathbf{x}\in\{0,1\}^{|V|}\subset\Delta^{|V|}$ as one-hot vectors, where $|V|$ is the vocabulary size and $\Delta^{|V|}$ is the simplex over this vocabulary. For finite-time processes, we let $T$ be the number of time steps. For each time step $t(i)=\frac{i}{T}\in[0,1]$ (i.e., for $i=0,1,\ldots,T$), each intermediate latent variable $\mathbf{z}_{t(i)}\in\{0,1\}^{|V|}$ has the marginal:
@@
Masked Diffusion Language Models  A special case of this formulation, known as ‘masking’ or ‘absorbing state’ diffusion, sets the limiting distribution $\bm{\pi}=\bm{m},$ a one-hot representation of a special [MASK] token. With $s(i)=\frac{i-1}{T}$ as the time step directly preceding $t$, these posteriors take the form:
@@
Importantly, despite the recent success of MDLM [38] and other absorbing state discrete diffusion models, these models suffer from a key drawback, which we call the failure to remask property. Namely, the posterior formula in (2) implies that any unmasked token $\mathbf{z}_{t}\neq\bm{m}$ must remain unchanged throughout the entire denoising process, for all $s\leq t,$ which Sahoo et al. [38] formalize as $q(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})=\mathrm{Cat}(\mathbf{z}_{s};\mathbf{z}_{t})$ if $\mathbf{z}_{t}\neq\bm{m}$. Once a token is decoded, this prediction is ‘locked-in’, a limitation that these diffusion models share with AR models.
As in Sahoo et al. [38], a denoising neural network $\mathbf{x}_{\theta}$ is used for parameterizing $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})=q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{\theta}(\mathbf{z}_{t}))$ and trained using the following objective:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}_{0}\sim q(\mathbf{z}_{0}\mid\mathbf{x})}\bigg{[}-\log p_{\theta}(\mathbf{x}\mid\mathbf{z}_{0})\bigg{]}+\mathbb{E}_{t\in\{\frac{1}{T},\ldots,1\}}\mathbb{E}_{\mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{x})}T\bigg{[}\frac{\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\log(\mathbf{x}_{\theta}(\mathbf{z}_{t})^{\top}\mathbf{x})\bigg{]}$ (3)
@@ 3 Remasking Diffusion Models
In this work, we alleviate the failure to remask limitation that affects state-of-the-art masked diffusion models. Our strategy is to design a probabilistic model similar to MDLM [38], but whose denoising posterior generalizes (2) and supports remasking.
Specifically, we will define a discrete diffusion model with posterior for $\mathbf{z}_{t}\neq\bm{m}$ given by:
@@ 3.1 ReMDM Forward Processes
where $q_{\sigma}(\mathbf{z}_{1}\mid\mathbf{x})=q(\mathbf{z}_{1})=\bm{m}$, as in masked diffusion. We construct the posteriors as:
$q_{\sigma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})=\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t})\mathbf{x}+\sigma_{t}\bm{m}),\quad\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}+\frac{1-\alpha_{s}-\sigma_{t}\alpha_{t}}{1-\alpha_{t}}\bm{m}),\quad\mathbf{z}_{t}=\bm{m}.\end{cases}$ (6)
Observe how (51) maintains the remasking property laid out in (4). Additionally, our chosen form for $q(\mathbf{z}_{s}\mid\mathbf{z}_{t}=\mathbf{m},\mathbf{x})$ ensures that the marginals for $q_{\sigma}(\mathbf{z}_{t}\mid\mathbf{x})$ are the same as in classical masked diffusion (e.g., Sahoo et al. [38]). This property will yield an ELBO for ReMDM that is very similar to that of MDLM, and will help us argue for using the ReMDM sampling procedure on top of a pretrained MDLM model. Formally, we establish the following result (proof in Appendix A.1):
@@
In Appendix A.2, we demonstrate that this ability to remask necessitates that ReMDM is a non-Markovian process (as in e.g., DDIM [48]). Intuitively, this is necessary because the latents $\mathbf{z}_{t}$ can transition from the masked state $\bm{m}$ back to $\mathbf{x}$, and therefore the transition kernel needs to be conditioned on $\mathbf{x}$.
@@ 3.2 Negative Evidence Lower Bound
$\mathcal{L}^{\sigma}=\mathbb{E}_{\mathbf{z}_{0}\sim q(\mathbf{z}_{0}\mid\mathbf{x})}\bigg{[}-\log p_{\theta}(\mathbf{x}\mid\mathbf{z}_{0})\bigg{]}+\mathbb{E}_{t\in\{\frac{1}{T},\ldots,1\}}\mathbb{E}_{\mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{x})}T\bigg{[}\frac{(1-\sigma_{t})\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\log(\mathbf{x}_{\theta}^{\top}\mathbf{x})\bigg{]}.$
Since $\alpha_{t}\leq\alpha_{s},$ each term in the second expectation increases monotonically in $\sigma_{t}$. When $\sigma_{t}=0$, for all $t$, we recover the MDLM objective from (3). This implies that the MDLM NELBO produces a tighter bound compared to ReMDM. Furthermore, since the diffusion loss term in (3.2) is simply a reweighted version of that in (3), one can presumably reuse weights from a model trained with (3), i.e., using different $\sigma_{t}$ at training and inference. Indeed, we find this to be a performant strategy in practice. Moreover, we find the performance (test perplexity) between models trained with the MDLM and the ReMDM objectives to be comparable (see Appendix E.1), further justifying our reuse of pretrained weights, with the added benefit of more flexible sampling unlocked by ReMDM.
@@ 4 ReMDM Samplers
In practice, we recommend using ReMDM with a different $\sigma_{t}$ at training and inference, just as one might switch to a different noise schedule during sampling. In particular, using $\sigma_{t}=0$ for training is equivalent to combining the ReMDM sampler with a pretrained MDLM model $\mathbf{x}_{\theta}$, which works well and does not require re-training $\mathbf{x}_{\theta}$.
@@ 4.1 Design Strategies for the Remasking Schedule $\sigma_{t}$
Here, we list several strategies for setting $\sigma_{t}\in[0,\sigma_{t}^{max}]$, which can either be employed separately or in conjunction with the strategies introduced in Section 4.2. For sequences, each token can have a different $\sigma_{t}^{(\ell)}$. When we omit this superscript, this implies that the same $\sigma_{t}$ is used for all tokens.
@@
Confidence-Based Schedule  In conjunction with the two strategies above, we explore a further reweighing of $\sigma_{t}$ which is based on the intuition that tokens of which the denoising model is less confident should be assigned a larger probability of remasking. Consider the $\ell$-th token in a sequence of $L$ latents at time $t$. For each $\ell\in\{1,\ldots,L\}$, we store its decoding probability at the time $\tau$ at which it was last unmasked. Concretely, if $\mathbf{z}_{t}^{(\ell)}\neq\bm{m}$, then we define $\psi_{t}^{(\ell)}:=\mathbf{x}_{\theta,\tau}^{(\ell)\top}\mathbf{z}_{\tau}^{(\ell)}$. If $\mathbf{z}_{t}=\bm{m}$, then $\psi_{t}^{(\ell)}:=\infty$. Thus, $\psi_{t}^{(\ell)}$ serves as a ‘confidence score’ for unmasked tokens. We then compute
@@
$p_{\theta}(\mathbf{z}_{s}\mid\mathbf{z}_{t})=q_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\sigma}}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}=\mathbf{x}_{\theta}(\mathbf{z}_{t}))={\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t})\mathbf{x}_{\theta}+\sigma_{t}\bm{m}),&\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}_{\theta}+\frac{1-\alpha_{s}-\sigma_{t}\alpha_{t}}{1-\alpha_{t}}\bm{m}),&\mathbf{z}_{t}=\bm{m}\\ \end{cases}}$